Reports 1-1 of 1 Clear search Modify search
Injection system (Laser power stabilization)
derossi, chiummo, cleva, coulon - 15:55 Monday 04 November 2019 (47534) Print this report
PSTAB oscillation at 12 kHz

Since Oct 23rd we observe some strange oscillations at 12 kHz on the power stabilization error signal (plots 1 and 2). Looking at the fast channels it seems triggered by a saturation of the corrections occurring at 3.5 V.

This started after having lowered the impinging power on the photodiodes and having increased the gain and it can be explained by some saturation occuring in a component of the locking electronics.

As a temporary solution we increased again the power and lowered the gain, as it used to be before Oct 22nd. The loop seems more robust, even if the rms on the error signal has increased (plot 3).

In the meanwhile J.P. Coulon will do some modifications on the spare locking electronics to overcome this issue.

 

 

Images attached to this report
Comments to this report:
chiummo, DeRossi - 16:38 Monday 04 November 2019 (47540) Print this report

The attached plot shows a qualitative improvement over the frequency and ampitude of the glitches as seen by PSTAB_PD2 after the increase of power on the photodiodes. Investigation in progress.

 

Images attached to this comment
Search Help
×

Warning

×